OpenStudy (erinweeks):

Help would be appreciated (: Find the values of the 30th and 90th percentiles of the data. 129, 113, 200, 100, 105, 132, 100, 176, 146, 152 A. 30th percentile = 105; 90th percentile = 200 B. 30th percentile = 113; 90th percentile = 200 C. 30th percentile = 105; 90th percentile = 176 D. 30th percentile = 113; 90th percentile = 176

OpenStudy (erinweeks):

@jim_thompson5910 .. do you know this ?

jimthompson5910 (jim_thompson5910):

First sort the values 129, 113, 200, 100, 105, 132, 100, 176, 146, 152 becomes 100, 100, 105, 113, 129, 132, 146, 152, 176, 200

jimthompson5910 (jim_thompson5910):

The 30th percentile is the point where 30% of the data is below that value

OpenStudy (erinweeks):

but them in order . okay ! i did that part . & what does that mean ?

jimthompson5910 (jim_thompson5910):

So if you have say 100 values, then 30% of that (0.3*100 = 30 values) are below the 30th percentile

OpenStudy (erinweeks):

so in this case we have 10 right .

jimthompson5910 (jim_thompson5910):

yes, I'm looking up a percentile formula now, one sec

OpenStudy (erinweeks):

okay !

jimthompson5910 (jim_thompson5910):

hmm there are multiple percentile formulas, so it depends on what your book says...do you have your book with you?

OpenStudy (erinweeks):

yes let me look quick !

OpenStudy (erinweeks):

my book doesnt have a formula .. it just says percentiles : seperate data sets into 100 equal parts

jimthompson5910 (jim_thompson5910):

well i found one formula to be n = (P/100)*N + 1/2 So the 30th percentile is at... n = (30/100)*10 + 1/2 n = 300/100 + 1/2 n = 3 + 1/2 n = 3.5 which rounds to n = 4, so the 4th value (which is 113) is the 30th percentile

jimthompson5910 (jim_thompson5910):

If we use the same formula for the 90th percentile, then... n = (P/100)*N + 1/2 So the 30th percentile is at... n = (90/100)*10 + 1/2 n = 900/100 + 1/2 n = 9 + 1/2 n = 9.5 which rounds to 10 so the 10th value, which is 200, is the 90th percentile

jimthompson5910 (jim_thompson5910):

so if we're using the correct formula (I hope we are), then the answer is B. 30th percentile = 113; 90th percentile = 200

OpenStudy (erinweeks):

okay i get that ! (: thank you so much

jimthompson5910 (jim_thompson5910):

let me know if it's the right answer or not

OpenStudy (anonymous):

yup, you were right :)

jimthompson5910 (jim_thompson5910):

thanks for the confirmation

OpenStudy (anonymous):

thanks for explaining :)